# Group-weighted conformal prediction

### v2

@@ 4.1 Theoretical guarantee
Finally, the quantity inside the last expected value must be $\geq 1-\alpha^{\prime}$, almost surely, by definition of the weighted empirical distribution $\widehat{P}_{\textnormal{score}}$ (see, e.g., Harrison, (2012, Lemma A.1)).
@@
$\mathbb{P}_{Q^{\prime}}\left\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\right\}\geq 1-\alpha^{\prime}-\min_{k:n_{k}>0}\{q^{\prime}_{k}/n_{k}\},$
@@
$\mathbb{P}_{Q}\left\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\right\}\geq q_{+}\cdot\left[1-\alpha^{\prime}-\min_{k:n_{k}>0}\{q^{\prime}_{k}/n_{k}\}\right]=1-\alpha-\min_{k:n_{k}>0}\{q_{k}/n_{k}\},$
@@ 5.1 Theoretical guarantee
The proof is given in Appendix A.1; essentially it is achieved by simply applying the guarantee for fixed $n_{k}$’s, given in Theorem 4.1, and then marginalizing over the random draw of the $n_{k}$’s.
@@ 5.3 Comparison to existing guarantees
We next provide an empirical comparison between the theoretical guarantee obtained in Corollary 5.2, which provides a lower bound for predictive coverage under group-wise covariate shift, against the existing guarantee obtained by Lei and Candès, (2021), shown above in (3), where the loss of coverage is bounded as $\mathbb{E}\left[\frac{\left|\widehat{w}(X)-w(X)\right|}{2}\right]$. We use the same settings as in Section 3.2, with three regimes: $K=10$ (a constant number of groups), $K=\lfloor\sqrt{n}\rfloor$ (a slowly growing number of groups), and $K=n/10$ (a proportional number of groups). The training and test distributions are again uniform, $p_{k}=q_{k}\equiv 1/K$. Here the coverage guarantee given by Lei and Candès, (2021) is the same as that shown in Figure 1, above; our new result, the coverage guarantee of Corollary 5.2, is now added to the plot for comparison. Note that this guarantee is for the version (12) of the group-weighted conformal prediction interval, which does not assume that we know the total number of groups.
Figure 3: Lei and Candès, (2021)’s WCP coverage guarantee (3), compared to the new guarantee given in Corollary 5.2, for varying numbers of groups. The horizontal line marks the target coverage level 90% (i.e., $\alpha=0.1$). See Section 5.3 for details.
As in Section 3.2, to ensure that $\widehat{w}$ is always well-defined even though there is a nonzero probability that not all groups are observed, we define $\widehat{w}(x)$ using estimated proportions $\widehat{p}_{k}=\frac{n_{k}+1}{n+K}$, i.e., augmenting each group count by 1. Therefore, we are giving two advantages to Lei and Candès, (2021)’s existing bound relative to ours: (1) the definition of $\widehat{p}_{k}$, which (since the true $p_{k}$’s are uniform) is more accurate than the actual empirical distribution; (2) Lei and Candès, (2021)’s coverage guarantee applies to the more conservative form of the method, given in (8), relative to ours.
@@ 6 Discussion
More broadly, this problem can be viewed as an instance of a more general class of settings, where covariate shift can be expressed via a finite dimensional parametrization. Whether these techniques can be extended to broader scenarios, moving beyond the setting of disjoint subpopulations, is an important direction for further exploration.
@@ A.3 Which data should be used to estimate the weights?
Recall that in Section 5.3, we showed an empirical comparison between our new bunds and the coverage guarantees established by Lei and Candès, (2021)’s existing result (3). In addition to the more rapid convergence of the lower bound in our new result, there is another interesting difference as compared to existing work. Specifically, existing theory (including Lei and Candès, (2021)’s result, but also related works such as Candès et al., (2023); Gui et al., (2022); Yin et al., (2022)) applies to an estimated weight function $\widehat{w}(x)$ that was defined independently of the calibration set. For example, if a pretraining data set was used for defining the score function (e.g., fitting a model $\widehat{f}:\mathcal{X}\rightarrow\mathbb{R}$, so that the score function is given by the residual score, $s(x,y)=|y-\widehat{f}(x)|$), the weight function $\widehat{w}(x)$ might also be trained ahead of time by estimating the covariate distribution $P_{X}$ using this pretraining data set. In contrast, the weights used in our proposed construction for the GWCP method specifically use the calibration set’s proportions—recall from Section 3.1 that we can think of GWCP as essentially running WCP with weight function $\widehat{w}(X)=q_{k}/(n_{k}/n)$ for $X^{0}=k$, where $n_{k}$ is the number of calibration data points in group $k$.
For both types of results, the particular choice is essential to the theory: Lei and Candès, (2021)’s result relies on $\widehat{w}$ being independent of the calibration set, while in contrast, our proofs rely on using the $n_{k}$’s observed in the calibration set. Which approach is the “right” one for achieving the best performance—that is, for the least loss of coverage?
@@
    $*.\widehat{w}_{k}=q_{k}/\widehat{p}_{k}\textnormal{ where }\widehat{p}_{k}=\frac{\sum_{i=1}^{n_{*}}\mathbf{1}_{X_{i}^{*}{}^{0}=k}}{n_{*}}.$
@@ A.4 A corrected GWCP for exact coverage
where (as in the proof of Theorem 4.1) the last step holds by definition of the weighted empirical distribution $\widehat{P}_{\textnormal{score}}$ (see, e.g., Harrison, (2012, Lemma A.1)). ∎

### v4

@@ 4.1 Theoretical guarantee
Finally, the quantity inside the last expected value must be $\geq 1-\alpha^{\prime}$, almost surely, by definition of the weighted empirical distribution $\widehat{P}_{\textnormal{score}}$ (see, e.g., Harrison (2012, Lemma A.1)).
@@
$\mathbb{P}_{Q^{\prime}}\left\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\right\}\geq 1-\alpha^{\prime}-\max_{k:n_{k}>0}\{q_{k}/n_{k}\},$
@@
$\mathbb{P}_{Q}\left\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\right\}\geq q_{+}\cdot\left[1-\alpha^{\prime}-\min_{k:n_{k}>0}\{q^{\prime}_{k}/n_{k}\}\right]=1-\alpha-\max_{k:n_{k}>0}\{q_{k}/n_{k}\},$
@@ 5.1 Theoretical guarantee
The proof is given in Appendix A.1; essentially it is achieved by simply applying the guarantee for fixed $n_{k}$’s, given in Theorem 4.1, and then marginalizing over the random draw of the $n_{k}$’s. The lower bound on $p_{k}$ of $O\left(\frac{\log n}{n}\right)$ is a mild condition that ensures that all groups are observed with high probability.
@@ 5.2 Uniform groups and unobserved groups: a closer look
5.3 Unknown test group probabilities
In this section, we consider the setting where the the group probabilities in the test distribution (i.e., $q_{1},q_{2},\ldots,q_{K}$) are not known. Instead we have $m$ many draws $\tilde{X}^{0}_{1},\ldots,\tilde{X}^{0}_{m}\in\{1,\dots,K\}$ that offer an empirical approximation to this distribution over groups—for instance, we might have a sample of size $m$ from the test population, and the $\tilde{X}^{0}_{i}$’s are the group assignments of these $m$ data points.
Can we simply use the empirical proportions in each group $k=1,\dots,K$ in place of the $q_{k}$’s when constructing the GWCP prediction set (6)? That is, our empirically weighted prediction set is given by
$\widehat{C}_{n}(x)=\left\{y\in\mathcal{Y}:s(x,y)\leq\widehat{q}\right\}\textnormal{ where }\widehat{q}=\textnormal{Quantile}_{1-\alpha}\left(\sum_{k=1}^{K}\tilde{q}_{k}\widehat{P}_{\textnormal{score}}^{(k)}\right),\ \tilde{q}_{k}=\frac{\sum_{i=1}^{m}\mathbf{1}\{\tilde{X}^{0}_{i}=k\}}{m}.$ (13)
The following result verifies that this approach offers nearly the same level of coverage as before—the additional loss of coverage in the guarantee is only $\frac{1}{m+1}$.
Corollary 5.3.
Suppose the training data points $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ are sampled i.i.d. from the distribution $P$ as defined in (4), and the test point is independently drawn as $(X_{n+1},Y_{n+1})\sim Q$ as defined in (5). Moreover, suppose $\tilde{X}^{0}_{1},\dots,\tilde{X}^{0}_{m}$ are i.i.d. draws from the Multinomial$(q_{1},\dots,q_{K})$ distribution, drawn independently of the data. Then, for any fixed (i.e., pretrained) score function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, the prediction interval $\widehat{C}_{n}$ defined in (13) satisfies
$\mathbb{P}\left\{Y_{n+m+1}\in\widehat{C}_{n}(X_{n+m+1})\right\}\geq 1-\alpha-\frac{1}{m+1}-\mathbb{E}\left[\max_{k:n_{k}>0}\{\tilde{q}_{k}/n_{k}\}\right].$
The proof is given in Appendix A.1.
@@ 5.4 Comparison to existing guarantees
We next provide an empirical comparison between the theoretical guarantee obtained in Corollary 5.2, which provides a lower bound for predictive coverage under group-wise covariate shift, against the existing guarantee obtained by (Lei and Candès, 2021), shown above in (3), where the loss of coverage is bounded as $\mathbb{E}\left[\frac{\left|\widehat{w}(X)-w(X)\right|}{2}\right]$. We use the same settings as in Section 3.2, with three regimes: $K=10$ (a constant number of groups), $K=\lfloor\sqrt{n}\rfloor$ (a slowly growing number of groups), and $K=n/10$ (a proportional number of groups). The training and test distributions are again uniform, $p_{k}=q_{k}\equiv 1/K$. Here the coverage guarantee given by (Lei and Candès, 2021) is the same as that shown in Figure 1, above; our new result, the coverage guarantee of Corollary 5.2, is now added to the plot for comparison. Note that this guarantee is for the version (12) of the group-weighted conformal prediction interval, which does not assume that we know the total number of groups.
Figure 3: (Lei and Candès, 2021)’s WCP coverage guarantee (3), compared to the new guarantee given in Corollary 5.2, for varying numbers of groups. The horizontal line marks the target coverage level 90% (i.e., $\alpha=0.1$). See Section 5.4 for details.
As in Section 3.2, to ensure that $\widehat{w}$ is always well-defined even though there is a nonzero probability that not all groups are observed, we define $\widehat{w}(x)$ using estimated proportions $\widehat{p}_{k}=\frac{n_{k}+1}{n+K}$, i.e., augmenting each group count by 1. Therefore, we are giving two advantages to (Lei and Candès, 2021)’s existing bound relative to ours: (1) the definition of $\widehat{p}_{k}$, which (since the true $p_{k}$’s are uniform) is more accurate than the actual empirical distribution; (2) (Lei and Candès, 2021)’s coverage guarantee applies to the more conservative form of the method, given in (8), relative to ours.
@@ 6 Discussion
More broadly, this problem can be viewed as an instance of a more general class of settings, where covariate shift can be expressed via a finite dimensional parametrization. Whether these techniques can be extended to broader scenarios, moving beyond the setting of disjoint subpopulations, is an important direction for further exploration. Another future direction of interest is the problem of allowing for noisy or corrupted group labels in the test or training data, to handle scenarios where group labels may not be known exactly—for instance, due to issues such as privacy constraints or inaccurate data collection.
@@ A.1 Additional proofs
Proof of Corollary 5.3.
First we consider an equivalent representation of the problem. Suppose we have $m+1$ test points $(X_{n+1},Y_{n+1}),(X_{n+2},Y_{n+2}),\dots,(X_{n+m+1},Y_{n+m+1})$ sampled i.i.d. from $Q$; we will examine the coverage guarantee for the first test point $(X_{n+1},Y_{n+1})$ as before, while taking the remaining $m$ test points as our source of information for estimating the $q_{k}$’s—that is, we will define
$\tilde{q}_{k}=\frac{\sum_{i=1}^{m}\mathbf{1}\{X^{0}_{n+1+i}=k\}}{m}.$
Now let $\tilde{Q}=\sum_{k=1}^{K}\tilde{q}_{k}\cdot\widehat{P}_{\text{score}}^{(k)}$, and let $\tilde{Q}^{*}=\sum_{k=1}^{K}\tilde{q}^{*}_{k}\cdot\widehat{P}_{\text{score}}^{(k)}$ where
$\tilde{q}^{*}_{k}=\frac{\sum_{i=1}^{m+1}\mathbf{1}\{X^{0}_{n+i}=k\}}{m+1}$
is the empirical frequency of group $k$ in the entire test set (i.e., the test point of interest, $(X_{n+1},Y_{n+1})$, along with the $m$ test points used to construct the $\tilde{q}_{k}$’s). Note that by definition we have
$\widehat{q}=\textnormal{Quantile}_{1-\alpha}(\tilde{Q}).$
Furthermore,
$\tilde{q}^{*}_{k}=\frac{m}{m+1}\tilde{q}_{k}+\frac{\mathbf{1}\{X^{0}_{n+1}=k\}}{m+1},$
and in particular, this means that we can calculate
$\textnormal{d}_{\text{TV}}\left(\tilde{Q},\tilde{Q}^{*}\right)=\frac{1}{2}\sum_{k=1}^{K}|\tilde{q}_{k}-\tilde{q}^{*}_{k}|=\frac{1}{2}\sum_{k=1}^{K}\frac{1}{m+1}\left|\tilde{q}_{k}-\mathbf{1}\{X^{0}_{n+1}=k\}\right|\leq\frac{1}{m+1}.$
Thus, $\widehat{q}=\textnormal{Quantile}_{1-\alpha}(\tilde{Q})\geq\textnormal{Quantile}_{1-\alpha^{\prime}}(\tilde{Q}^{*})$ where $\alpha^{\prime}=\alpha+\frac{1}{m+1}$, which further implies that
$\widehat{C}_{n}^{*}(x):=\left\{y\in\mathcal{Y}:s(x,y)\leq\textnormal{Quantile}_{1-\alpha^{\prime}}(\tilde{Q}^{*})\right\}\subseteq\widehat{C}_{n}(x)$
. Therefore,
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\right\}\geq\mathbb{P}\left\{Y_{n+1}\in\widehat{C}_{n}^{*}(X_{n+1})\right\}.$
But by symmetry, since the $\tilde{q}_{k}$’s are constructed as a symmetric function of the $m+1$ test points, we have
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}_{n}^{*}(X_{n+1})\right\}=\mathbb{P}\left\{Y_{n+i}\in\widehat{C}_{n}^{*}(X_{n+i})\right\}$
for all $i=1,\dots,m+1$. Taking the average, then,
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}_{n}^{*}(X_{n+1})\right\}=\frac{1}{m+1}\sum_{i=1}^{m+1}\mathbb{P}\left\{Y_{n+i}\in\widehat{C}_{n}^{*}(X_{n+i})\right\}=\mathbb{E}\left[\frac{1}{m+1}\sum_{i=1}^{m+1}\mathbf{1}\left\{Y_{n+i}\in\widehat{C}_{n}^{*}(X_{n+i})\right\}\right].$
By the tower law we can write this as
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}_{n}^{*}(X_{n+1})\right\}=\mathbb{E}\left[\mathbb{E}\left[\frac{1}{m+1}\sum_{i=1}^{m+1}\mathbf{1}\left\{Y_{n+i}\in\widehat{C}_{n}^{*}(X_{n+i})\right\}\,\middle|\,X^{0}_{n+1},\dots,X^{0}_{n+m+1}\right]\right].$
Equivalently, sampling an index $I$ uniformly from $\{n+1,\dots,n+m+1\}$,
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}_{n}^{*}(X_{n+1})\right\}=\mathbb{E}\left[\mathbb{P}\left\{Y_{n+I}\in\widehat{C}_{n}^{*}(X_{n+I})\,\middle|\,X^{0}_{n+1},\dots,X^{0}_{n+m+1}\right\}\right].$
But conditional on the group assignments $X^{0}_{n+1},\dots,X^{0}_{n+m+1}$, we can observe that $(X_{n+I},Y_{n+I})$ is a draw from the distribution $\tilde{Q}^{*}$. Applying Theorem 5.1 (with $\tilde{Q}^{*}$ in place of $Q$), then,
$\mathbb{P}\left\{Y_{n+I}\in\widehat{C}_{n}^{*}(X_{n+I})\,\middle|\,X^{0}_{n+1},\dots,X^{0}_{n+m+1}\right\}\geq 1-\alpha^{\prime}-\max_{k:n_{k}>0}\{\tilde{q}^{*}_{k}/n_{k}\}$
and so after marginalizing,
$\mathbb{P}\left\{Y_{n+1}\in\widehat{C}_{n}^{*}(X_{n+1})\right\}\geq 1-\alpha^{\prime}-\mathbb{E}\left[\max_{k:n_{k}>0}\{\tilde{q}^{*}_{k}/n_{k}\}\right].$
Finally, we have
$\mathbb{E}\left[(\tilde{q}_{1},\dots,\tilde{q}_{K})\mid(\tilde{q}^{*}_{1},\dots,\tilde{q}^{*}_{K}),(n_{1},\dots,n_{K})\right]=(\tilde{q}^{*}_{1},\dots,\tilde{q}^{*}_{K}),$
so by Jensen’s inequality,
$\mathbb{E}\left[\max_{k:n_{k}>0}\{\tilde{q}_{k}/n_{k}\}\,\middle|\,(\tilde{q}^{*}_{1},\dots,\tilde{q}^{*}_{K}),(n_{1},\dots,n_{K})\right]\geq\max_{k:n_{k}>0}\{\tilde{q}^{*}_{k}/n_{k}\}.$
This completes the proof. ∎
@@ A.3 Which data should be used to estimate the weights?
Recall that in Section 5.4, we showed an empirical comparison between our new bounds and the coverage guarantees established by [Lei and Candès, 2021]’s existing result (3). In addition to the more rapid convergence of the lower bound in our new result, there is another interesting difference as compared to existing work. Specifically, existing theory (including [Lei and Candès, 2021]’s result, but also related works such as [Candès et al., 2023, Gui et al., 2022, Yin et al., 2022]) applies to an estimated weight function $\widehat{w}(x)$ that was defined independently of the calibration set. For example, if a pretraining data set was used for defining the score function (e.g., fitting a model $\widehat{f}:\mathcal{X}\rightarrow\mathbb{R}$, so that the score function is given by the residual score, $s(x,y)=|y-\widehat{f}(x)|$), the weight function $\widehat{w}(x)$ might also be trained ahead of time by estimating the covariate distribution $P_{X}$ using this pretraining data set. In contrast, the weights used in our proposed construction for the GWCP method specifically use the calibration set’s proportions—recall from Section 3.1 that we can think of GWCP as essentially running WCP with weight function $\widehat{w}(X)=q_{k}/(n_{k}/n)$ for $X^{0}=k$, where $n_{k}$ is the number of calibration data points in group $k$.
For both types of results, the particular choice is essential to the theory: [Lei and Candès, 2021]’s result relies on $\widehat{w}$ being independent of the calibration set, while in contrast, our proofs rely on using the $n_{k}$’s observed in the calibration set. Which approach is the “right” one for achieving the best performance—that is, for the least loss of coverage?
@@
    $\widehat{w}_{k}=q_{k}/\widehat{p}_{k}\textnormal{ where }\widehat{p}_{k}=\frac{\sum_{i=1}^{n_{*}}\mathbf{1}_{X_{i}^{*}{}^{0}=k}}{n_{*}}.$
@@ A.4 A corrected GWCP for exact coverage
where (as in the proof of Theorem 4.1) the last step holds by definition of the weighted empirical distribution $\widehat{P}_{\textnormal{score}}$ (see, e.g., Harrison [2012, Lemma A.1]). ∎